# Voting by mail: a Markov chain model for managing the security risks of absentee voting

### v2

@@ 3.4 Mitigations Layer
The mitigation layer connects with both the process and attack layers. We model structural mitigations both the physical pathways (arcs) and the probabilities (transitions) of a ballot’s progression. These require specific responsive measures from voters to negate the effects of attacks. The DTMC model consists of five structural mitigations ; Table 2 details each mitigation. We derived these mitigations from prior work that condense countermeasures published by CISA into practical logic intertwined with the VBM process [9, 5]. Table 2 presents these mitigations, categorized according to their operational nature along with the controlling entity, election office or voter.
@@ 4 DTMC Model
In this section, we introduce a DTMC model of the VBM system based on a multi-layer configuration of the process, attacks, and mitigations layers. The DTMC model captures the stochastic movement of ballots through various stages from a ballot request to the counting of ballots. This approach allows us to analyze how different factors influence overall election performance. A crucial aspect of the DTMC model is its ability to delineate the interaction between the terminal attack nodes, the mitigation layer, and the VBM process.
@@
On election day, there are transition probability arcs re-positioned to move a ballot to its final post-election status in the last period, , to evaluate the performance measures. Given a random process with finite states, the one-step transition probability of the process moving from state in time step to state in a single time step is
We define a transition probability matrix for each time step () to calculate the ballot states at the end of each time step [28]. Figure 2 illustrates the DTMC model state diagram for , showing the final ballot states along with three distinct layers: the process layer, the attack layer, and the mitigation layer. The DTMC model comprises 30 finite states , categorized into six recurrent final ballot states and 24 transient states. Next, we describe the DTMC states and transition probabilities, starting with the process layer and then adding the attack and mitigation layers.
@@
Additionally, Figure 2 uses line styles to convey information about the arc transition probabilities in the DTMC. Solid lines indicate transitions with non-zero probabilities at all steps, except at the final time step . In contrast, dotted and dashed lines represent transitions with non-zero probabilities only at specific time steps, such as those associated with transient malicious attacks. The complexity of the model leads to overlapping lines in the figure; intersections marked by a dot signal a connection to the intersecting line. In the model, masquerade attacks change a ballot’s status and add a hidden attribute. Process layer states are mirrored in the attacks layer and labeled with an ‘A’ in Figure 2 to indicate the altered status of these ballots.
@@
At the end of the time horizon, the ballots move to the final ballot states. Figure 3 shows the non-zero transition probabilities on election day at time step . If ballots are not returned to the election office, they are “not counted, not returned” (NC, NR). These six final ballot states are recurrent in the DTMC, and all others are transient. If ballots are returned to the election office at they are “not returned, late”(NC, L).
Next, we summarize the transitions. Let capture the transition probability matrix immediately after the time steps . Voters can request ballots at different times. Let capture the number of ballots requested at time , which reflects the distribution of times when voters request absentee ballots. For the ballots requested at time , let be the vector of the probability mass function for the starting states of the ballot, where all the ballots are initiated in state , that is, and for all other states . For a ballot requested at time , we can compute the vector of state probabilities at the end of the time horizon at time step after all ballots have been evaluated and transitioned to one of the ballot status states:
Then, we can compute the overall distribution of final ballot statuses as that reflects the expected number of ballots in each final ballot state.
@@ 5 Wisconsin 2020 General Election Case Study
We conduct a case study based on data from the 2020 General Election in Milwaukee County, Wisconsin. In 2020 Milwaukee County, the largest county in Wisconsin, had 939,489 residents and 478 voting wards [33, 4, 15]. Milwaukee County election officials used multiple absentee voting mechanisms during the 2020 General Election, including VBM with designated drop boxes and in-person absentee voting. The state of Wisconsin allowed absentee voting with “no excuses” and sent voter reminders to return ballots [17]. At that time, the county had 550,132 registered voters and experienced a turnout of 83.67 percent [38]. There are no known malicious attacks on VBM in Milwaukee County in the 2020 General Election, although there were two convictions elsewhere in Wisconsin associated with independent instances of VBM election fraud, each affecting a single ballot [32].
@@
Using this data, we determined the daily number of absentee ballots requested, , for , with 94.1 percent of these ballots being returned. The procedures and recourse actions available to voters during the General Election on November 3, 2020 are represented in Table 2 as mitigations. All requested ballots correspond to the starting state of the DTMC state . Additional data were obtained from the MIT Election Data + Science Lab and the United States Office of the Inspector General [16, 19]. The case study begins with establishing transition probabilities for each arc in the process for every time interval in the DTMC model.
@@ 5.1 Time
In the VBM process, ballot states of the DTMC model transitions on a daily basis. The case study begins on September 17, 2020 (), the first day ballots are mailed to voters, and ends after election day, November 4, 2020 (). The DTMC transition probabilities are time-dependent in accounting for procedure changes over the VBM timeline. For example, some mitigations are not available at all times due to VBM procedure, and malicious attacks occur at various times. However, many of the transition probabilities are homogeneous across time intervals. As a result, we partition the time horizon into four segments with transition probabilities that are time-homogeneous within an interval.
@@
5.2 Transition probabilities
We define the transition probability matrices as follows. We first define the transition probabilities for the baseline model that only considers non-malicious attacks. Non-malicious attacks are represented as consistent threats with time-inhomogenous probabilities of occurrence. Later, we consider malicious attacks with a duration of one day.
Table 4 outlines the transition probabilities for each arc active in the baseline model. The first and second columns define the two DTMC states associated with a non-zero transition probability. The next column provides a brief functional description of the ballot state transition that the arc represents. The columns labeled Interval 1, 2 and 3 represent the values of associated transition probability across time intervals 1, 2, and 3. We omit interval 4, since these transition probabilities are or with arcs illustrated in Figure 3. The justification column reports a reference for each transition probability value and denotes which arcs were adjusted through calibration. The arc configuration aligns with the model shown in Figure 2.
We discuss several transition probabilities. The transition probability from process state to is , which reflects the proportion of election mail that is processed on-time as reported by the USPS Inspector General [20]. The rest of the ballots remain in state . The transition from state to reflects the return of completed ballots by voters (that is, the voter rate of ballot return). This transition captures the proportion of voters who return the marked ballot within a day of receiving it in the mail, which is determined as a part of the model calibration since this value is not directly recorded. MIT Election Data + Science Lab [16] provide the values for utilization of drop boxes over USPS to return ballots. Transitions from state to state , which occur with a probability of , represents the proportion of ballots returned by drop boxes. [16] Transitions at the end of the time interval move ballots to their final ballot states with probability 1.0. These probabilities can be gleaned from Figure 3. All other transition probabilities are zero. We calibrated the DTMC using historical data from the WEC 2020 General Election report [40]. This involved adjusting the voter rate of ballot return, the arc probability from node to , to align the output of (C,U) and (NC,U) ballots with the observed ballot return rate seen in aggregate by the state of Wisconsin. Following calibration, we validate the model’s predictive accuracy by testing it against independent data validation points for Milwaukee County. This ensures blind prediction input data such that our model fits historical data and ensures the model is capable of making accurate predictions in varying scenarios.
@@ 5.3 Calibration
We calibrated the model using Wisconsin state-level data to align the recorded values of the ballots returned on various days with the calculated number of ballots counted and unaltered () final ballot state of the DTMC model. To accomplish this, we set the voter rate of ballot return to reflect the influence of mitigation (reminders); the arc transition probability from node to node reflects the voter rate of ballots to return in intervals 2 and 3.
@@
We then consider the mitigations of the model to determine a starting point for policies active in the 2020 Wisconsin General Election. We study the availability of mitigations M3, M4, M5, M6, and M7. Mitigation strength refers to the effect of each mitigation on the magnitude of the associated arc transition probabilities. Table 5 lists the inputs associated with mitigation strength in the baseline model. Low strength (less than 0.05) is associated with a minimal impact of that mitigation on the VBM system. Wisconsin voters were able to view ballot notifications manually through myvote.wi.gov, Wisconsin did not implement automatic ballot notification in 2020, so this value is low to reflect that few voters manually checked their ballot statuses. We set mitigation strength to 0.0265, which resulted in a mean absolute deviation of 0.24-percent from the observed number of (NC,U) ballots. Next, we set mitigation to 0.90 to represent the ability of voters to request replacement ballots. We set mitigation to 0.74 to represent the sent and advertised reminders to return ballots in the late second interval. In Table 4 we see the arc from node to node increases in interval 2 to account for the increased rate of ballot return caused by implementation of mitigation . Next, we set mitigation to 0.40 to represent the rate voters choose to submit an absentee ballot in person. Finally, we set the strength of to the average use in the state of Wisconsin [16]. This results in a mean absolute deviation of 0.25-percent from the observed number of (NC,U) ballots.
@@
The calibration of the baseline DTMC model also requires consideration of non-malicious attacks; we model as part of normal system operations. The ballot rejection rate of the model is associated with the arc transition probability from node to . It correlates directly with the number of ballots rejected by the election office due to voter error. We set the ballot rejection rate to 0.000162 to closely match the observed number of returned ballots that were unaltered and not counted (NC,U). Next, we vary the impact of the remaining non-malicious attacks, representing the ballots that never returned to the election office. Recall terminal attack node represents the accidental loss of ballots in the mail. The terminal attack node is the misaddressing of ballots to the voter, and represents the accidental loss of ballots in the election office. These non-malicious attacks all result in the same penalty, and the ballots impacted are considered not counted and not returned (NC,NR). Unlike the ballot rejection rate of the election office, there is a lack of evidence to support the different magnitudes for the non-malicious attacks . These non-malicious attack strength values are equivalent for all periods [37]. Therefore, we set the terminal attacks to equal strength. As a result, the following arcs have the same transition probability of 0.0343: arc from to ; arc from to ; arc from to . These arcs result in a mean absolute deviation of 0.14-percent from the observed (C,U) ballots.
@@ 5.4 Validation
Table 6 summarizes the validation points and relates them to the final ballot states of the model. The last two columns of Table 6 show the number of ballots under each validation point, comparing real world ballot counts “WEC ballot counts scaled to Milwaukee” and the DTMC model ballot counts. The validation process, aligning the DTMC model with real-world data from the WEC report, reinforces the model’s reliability in reflecting absentee voting behaviors and provides valuable insights into vote-by-mail dynamics in Wisconsin.The last column of Table 6 shows the DTMC model output values for the given validation points. Comparative analysis aligned the model’s final ballot states with the WEC report’s metrics. Once calibrated and validated the model, most transition probabilities were established for all time intervals. Milwaukee County data points are then used to establish a baseline model that can be used to consider multiple scenarios.
@@ 6 Computational Results
This section discusses the baseline DTMC model results for the Milwaukee County case study. The model was coded and implemented in Matlab, and the code and data is publicly available [8]. We then study the impact of several types of malicious attacks on the established baseline. We report the expected number of ballots in each of the final ballot states—rounded to the nearest ballot—under multiple scenarios.
@@ 6.1 Baseline model
We first analyze the VBM system under normal conditions with only non-malicious attacks in the baseline model introduced in Section 5. Non-malicious attacks (see Table 4) are included in the baseline model, since they are an uncontrolled, continuous part of the VBM process. Additionally, we perform a sensitivity analysis of four of the non-malicious attacks by varying their corresponding probabilities of occurring.
Table 7 summarizes the expected number of ballots in each final ballot state: counted and unaltered (C,U), not counted and unaltered (NC,U), correctly counted but altered (C,A), not counted due to late arrival (NC,L), not counted due to not being returned (NC,NR), and not counted and altered (NC,A). Table 7 also reports a sensitivity analysis that varies the election office rejection rate and three other non-malicious attack transition probabilities used in the model. The election office rejection rate varies based on ballots returned and rejected by the election office on election day due to non-malicious attacks and , according to subject matter expert testimony [1]. The non-malicious attack strength of attacks and represent ballots not returned due to either not being delivered to the voter or being discarded at the election office before counting the valid returned ballots on election day [40].
@@
The top row of Table 7 represents the baseline model, and it reports the expected number of ballots in each final ballot state. Each subsequent row represents a different scenario with varying attack parameters. The incremental increase of the rejection rate associated with the arc transition probability from node to by a probability of 0.00001 increases the number of (NC,U) ballots by 6-percent (35 total ballots), suggesting that efforts to reduce the rejection rate can decrease the number of ballots that are not counted. Higher attack strengths associated with increase the number of ballots not returned (NC,NR). For example, when the attach strength increases from to , the number of (NC,NR) ballots increase by . Additionally, we note that the election office rejection rate () for Milwaukee County is substantially lower than that of the other non-malicious attacks. These insights underscore the need for additional attention to unreturned (NC,NR) ballots in the VBM process and maintaining minimal non-malicious attack strengths to ensure that ballots are returned and counted.

### v3

@@ 3.4 Mitigations Layer
The mitigation layer connects with both the process and attack layers. We model structural mitigations in both the physical pathways (arcs) and the transitions of a ballot’s progression. In the context of the analysis, these transition values may represent relative likelihoods or stress-test intensities rather than empirically derived probabilities, allowing us to evaluate potential risk exposure in the absence of historical data. Table 2 details the five structural mitigation mitigations () that require specific responsive measures from voters to negate the effects of attacks. These mitigations derive from prior work that condense countermeasures published by CISA into practical logic intertwined with the VBM process [9, 36], and the mitigation numbering starts at 3 to be consistent with the original labels in [9]. Other mitigations exist [8, 9], however, we focus on the core set of mitigations that are in use during the voting process. Table 2 presents the mitigations categorized according to their operational nature along with the controlling entity, election office or voter.
@@
3.5 Modeling Assumptions
To ensure transparency in the application of the DTMC framework, we consolidate several key assumptions here. First, the model assumes that the high-level VBM process remains functionally consistent across different US municipalities. Second, the transition values used in the stress-test scenarios considered later in this paper represent relative likelihoods of risk exposure rather than strictly empirical probabilities, reflecting the inherent difficulty in estimating the frequency of rare security incidents. Third, the model treats the final ballot states (e.g., Counted, Unaltered (C,U)) as mutually exclusive and absorbing, allowing for the quantification of the final outcome distribution. Finally, while non-malicious risks are modeled as persistent baseline threats, malicious attacks are assumed to have a localized duration of typically one day to assess the impact to targeted attacks and identify specific times of heightened vulnerability within the election cycle.
@@ 4 DTMC Model
In this section, we introduce a DTMC model of the VBM system based on a multi-layer configuration of the process, attacks, and mitigations layers. The DTMC model captures the stochastic movement of ballots through various stages from a ballot request to the counting of ballots. It aggregates granular operational voting steps into functional steps to analyze how different factors influence overall election performance. A crucial aspect of the DTMC model is its ability to delineate the interaction between the terminal attack nodes, the mitigation layer, and the VBM process.
@@
On election day, the arcs are re-positioned to move a ballot to its final post-election status in the last period, , to evaluate the performance measures. Given a random process with finite states, the one-step transition probability of the process moving from state in time step to state in a single time step is
We define a state transition matrix for each time step () to calculate the ballot states at the end of each time step [29]. Figure 2 illustrates the DTMC model state diagram for , showing the final ballot states along with three distinct layers: the process layer, the attack layer, and the mitigation layer. The DTMC model comprises 30 finite states , categorized into six recurrent final ballot states and 24 transient states. Next, we describe the DTMC states and transitions, starting with the process layer and then adding the attack and mitigation layers.
@@
Additionally, Figure 2 uses line styles to convey information about the arc transitions in the DTMC. Solid lines indicate transitions with non-zero probabilities at all steps, except at the final time step . In contrast, dotted and dashed lines represent transitions with non-zero relative likelihoods only at specific time steps, such as those associated with transient malicious attacks. The complexity of the model leads to overlapping lines in the figure; intersections marked by a dot signal a connection to the intersecting line. In the model, masquerade attacks change a ballot’s status and add a hidden attribute. Process layer states are mirrored in the attacks layer and labeled with an ‘A’ in Figure 2 to indicate the altered status of these ballots.
@@
At the end of the time horizon, the ballots move to the final ballot states. Figure 3 shows the non-zero transition relative likelihoods on election day at time step . If ballots are not returned to the election office, they are “not counted, not returned” (NC, NR). These six final ballot states are recurrent in the DTMC, and all others are transient. If ballots are returned to the election office at they are “not returned, late”(NC, L).
Next, we summarize the transitions. Let capture the state transition matrix immediately after the time steps . Voters can request ballots at different times. Let capture the number of ballots requested at time , which reflects the distribution of times when voters request absentee ballots. For the ballots requested at time , let be the vector of the probability mass function for the starting states of the ballot, where all the ballots are initiated in state , that is, and for all other states . For a ballot requested at time , we can compute the vector of state probabilities at the end of the time horizon at time step after all ballots have been evaluated and transitioned to one of the ballot status states:
Then, we can compute the overall distribution of final ballot statuses as that reflects the expected number of ballots in each final ballot state. Note that several U.S. states send ballots to all registered voters. The DTMC model can be adapted to this situation by setting the initial ballot request parameter at time 1 () to the total number of registered voters and setting for .
@@ 5 Wisconsin 2020 General Election Case Study
We conduct a case study based on data from the 2020 General Election in Milwaukee County, Wisconsin. In 2020 Milwaukee County, the largest county in Wisconsin, had 939,489 residents and 478 voting wards [35, 4, 17]. Milwaukee County election officials used multiple absentee voting mechanisms during the 2020 General Election, including VBM with designated drop boxes and in-person absentee voting. The state of Wisconsin allowed absentee voting with “no excuses” (i.e., all voters are eligible for requesting an absentee ballot) and sent voter reminders to return ballots [19]. At that time, the county had 550,132 registered voters and experienced a turnout of 83.67 percent [40]. There are no known malicious attacks on VBM in Milwaukee County in the 2020 General Election, although there were two convictions elsewhere in Wisconsin associated with independent instances of VBM election fraud, each affecting a single ballot [34]. Certain low-volume attacks, such as the theft of a single blank ballot from a mailbox, are inherently difficult to detect and may be indistinguishable from benign process errors (e.g., accidental postal loss). As a result, we consider hypothetical worst-case threat scenarios when modeling the malicious attacks to provide insight into the defensive measures could reduce risk or threat exposure.
@@
Using this data, we determined the daily number of absentee ballots requested, , for , with 94.1 percent of these ballots being returned. The procedures and recourse actions available to voters during the General Election on November 3, 2020 are represented in Table 2 as mitigations. All requested ballots correspond to the starting state of the DTMC state . Additional data were obtained from the MIT Election Data + Science Lab and the United States Office of the Inspector General [18, 21]. The case study begins with establishing transition likelihoods for each arc in the process for every time interval in the DTMC model.
@@ 5.1 Time
In the VBM process, ballot states of the DTMC model transitions on a daily basis. The case study begins on September 17, 2020 (), the first day ballots are mailed to voters, and ends after election day, November 4, 2020 (). The DTMC transition likelihoods are time-dependent in accounting for procedure changes over the VBM timeline. For example, some mitigations are not available at all times due to VBM procedure, and malicious attacks occur at various times. However, many of the transition likelihoods are homogeneous across time intervals. As a result, we partition the time horizon into four segments with transitions that are time-homogeneous within an interval, not including those associated with malicious attacks that are described in Section 6.
@@
5.2 Transitions
We define the state transition matrices as follows. We first define the transitions for the baseline model that only considers non-malicious attacks. Non-malicious attacks are represented as consistent threats with time-inhomogenous relative likelihoods of occurrence. Later, we consider malicious attacks with a duration of one day.
Table 4 outlines the transition relative likelihoods for each arc active in the baseline model. The first and second columns define the two DTMC states associated with a non-zero transition relative likelihood. The next column provides a brief functional description of the ballot state transition that the arc represents. The columns labeled Interval 1, 2 and 3 represent the values of associated transition relative likelihood across time intervals 1, 2, and 3. We omit interval 4, since these transition relative likelihoods are or with arcs illustrated in Figure 3. The justification column reports a reference for each transition value and denotes which arcs were adjusted through calibration. The arc configuration aligns with the model shown in Figure 2.
We discuss several transitions. The transition relative likelihood from process state to is , which reflects the proportion of election mail that is processed on-time as reported by the USPS Inspector General [21]. The rest of the ballots remain in state . The transition from state to reflects the return of completed ballots by voters (that is, the voter rate of ballot return). This transition captures the proportion of voters who return the marked ballot within a day of receiving it in the mail, which is determined as a part of the model calibration since this value is not directly recorded. MIT Election Data + Science Lab [18] provide the values for utilization of drop boxes over USPS to return ballots. Transitions from state to state , which occur with a relative likelihood of , represents the proportion of ballots returned by drop boxes [18]. Transitions at the end of the time interval move ballots to their final ballot states with relative likelihood 1.0. These probabilities can be gleaned from Figure 3. All other transition relative likelihoods are zero. We calibrated the DTMC using historical data from the WEC 2020 General Election report [45]. This involved adjusting the voter rate of ballot return, the arc transition likelihood from node to , to align the output of (C,U) and (NC,U) ballots with the observed ballot return rate seen in aggregate by the state of Wisconsin. Following calibration, we validate the model’s predictive accuracy by testing it against independent data validation points for Milwaukee County. This ensures blind prediction input data such that our model fits historical data and ensures the model is capable of making accurate predictions in varying scenarios.
@@ 5.3 Calibration
We calibrated the model using Wisconsin state-level data to align the recorded empirical values of the ballots returned on various days with the calculated number of ballots counted and unaltered () final ballot state of the DTMC model. To accomplish this, we set the voter rate of ballot return to reflect the influence of mitigation (reminders); the arc transition from node to node reflects the voter rate of ballots to return in intervals 2 and 3.
@@
We then consider the mitigations of the model to determine a starting point for policies active in the 2020 Wisconsin General Election. We study the availability of mitigations M3, M4, M5, M6, and M7. Mitigation strength refers to the effect of each mitigation on the magnitude of the associated arc transitions. Therefore, the mitigation strength values are relative likelihood values. Table 5 lists the inputs associated with mitigation strength in the baseline model. Wisconsin did not implement automatic ballot notification in 2020, when voters were able to manually view ballot notifications through myvote.wi.gov. We set mitigation strength to 0.0265 to align the model’s baseline rejection rate with the empirical value of (NC,U) ballots. Next, we set mitigation to 0.90 to represent the ability of voters to request replacement ballots. We set mitigation to 0.74 to represent the sent and advertised reminders to return ballots in the late second interval. In Table 4 we see the arc from node to node increases in interval 2 to account for the increased rate of ballot return caused by implementation of mitigation . Next, we set mitigation to 0.40 to represent the rate voters choose to submit an absentee ballot in person. Finally, we set the strength of to the average use in the state of Wisconsin [18]. This results in a mean absolute deviation of 0.25-percent from the observed number of (NC,U) ballots.
@@
The calibration of the baseline DTMC model also requires consideration of non-malicious attacks; we model as part of normal system operations. The ballot rejection rate of the model is associated with the arc transition relative likelihood from node to . It correlates directly with the number of ballots rejected by the election office due to voter error. We set the ballot rejection rate to 0.000162 to closely match the observed number of returned ballots that were unaltered and not counted (NC,U). Next, we vary the impact of the remaining non-malicious attacks, representing the ballots that never returned to the election office. Recall terminal attack node represents the accidental loss of ballots in the mail. The terminal attack node is the misaddressing of ballots to the voter, and represents the accidental loss of ballots in the election office. These non-malicious attacks all result in the same penalty, and the ballots impacted are considered not counted and not returned (NC,NR). Unlike the ballot rejection rate of the election office, there is a lack of evidence to support the different magnitudes for the non-malicious attacks . These non-malicious attack strength values are equivalent for all periods [44]. Therefore, we set the terminal attacks to equal strength. As a result, the following arcs have the same transition relative likelihood of 0.0343: arc from to ; arc from to ; arc from to . These arcs result in a mean absolute deviation of 0.14-percent from the observed (C,U) ballots.
@@ 5.4 Validation
Table 6 summarizes the validation points and relates them to the final ballot states of the model. The last two columns of Table 6 show the number of ballots under each validation point, comparing real world ballot counts “WEC ballot counts scaled to Milwaukee” and the DTMC model ballot counts. The validation process, aligning the DTMC model with real-world data from the WEC report, reinforces the model’s reliability in reflecting absentee voting behaviors and provides valuable insights into vote-by-mail dynamics in Wisconsin.The last column of Table 6 shows the DTMC model output values for the given validation points. Comparative analysis aligned the model’s final ballot states with the WEC report’s metrics. Once calibrated and validated the model, most transitions were established for all time intervals. Milwaukee County data points are then used to establish a baseline model that can be used to consider multiple scenarios.
@@ 6 Computational Results
This section discusses the baseline DTMC model results for the Milwaukee County case study. The model is coded and implemented in Matlab, and the code and data are publicly available [10]. We then study the impact of several types of malicious attacks on the established baseline. We report the expected number of ballots in each of the final ballot states—rounded to the nearest ballot—under multiple scenarios.
@@ 6.1 Baseline model
We first analyze the VBM system under normal conditions with only non-malicious attacks in the baseline model introduced in Section 5. Non-malicious attacks (see Table 4) are included in the baseline model, since they are an uncontrolled, continuous part of the VBM process. Additionally, we perform a sensitivity analysis of four of the non-malicious attacks by varying their corresponding relative likelihoods of occurring.
Table 7 summarizes the expected number of ballots in each final ballot state: counted and unaltered (C,U), not counted and unaltered (NC,U), correctly counted but altered (C,A), not counted due to late arrival (NC,L), not counted due to not being returned (NC,NR), and not counted and altered (NC,A). Table 7 also reports a sensitivity analysis that varies the election office rejection rate and three other non-malicious attack transitions used in the model. The election office rejection rate varies based on ballots returned and rejected by the election office on election day due to non-malicious attacks and , according to subject matter expert testimony [1]. The non-malicious attack strength of attacks and represent ballots not returned due to either not being delivered to the voter or being discarded at the election office before counting the valid returned ballots on election day [45].
@@
The top row of Table 7 represents the baseline model, and it reports the expected number of ballots in each final ballot state. Each subsequent row represents a different scenario with varying attack parameters. The incremental increase of the rejection rate associated with the arc transition from node to by a value of 0.00001 increases the number of (NC,U) ballots by 6-percent (35 total ballots), suggesting that efforts to reduce the rejection rate can decrease the number of ballots that are not counted. Higher attack strengths associated with increase the number of ballots not returned (NC,NR). For example, when the attach strength increases from to , the number of (NC,NR) ballots increase by . Additionally, we note that the election office rejection rate () for Milwaukee County is substantially lower than that of the other non-malicious attacks. These insights underscore the need for additional attention to unreturned (NC,NR) ballots in the VBM process and maintaining minimal non-malicious attack strengths to ensure that ballots are returned and counted.